# MM-Prompt: Cross-Modal Prompt Tuning for Continual Visual Question Answering

### v1

@@ 1 Introduction
Pre-trained models (PTMs) have achieved considerable success in Visual Question Answering (VQA) [1, 2, 3], a traditional multi-modal task that generates answers to questions based on relevant images. Recently, researchers further extended PTMs to Continual VQA (CVQA) [4, 5, 6, 7, 8, 9], where new content emerges over time, with approaches developed to address the challenge of catastrophic forgetting. Prompt tuning, a PTM-based method proven effective in CL [10, 11, 12, 13, 14, 15, 16], which first selects prompts based on input representations via similarity, then injects the selected prompts into encoders to guide learning. This approach has shown good performance in CVQA [6, 4, 8, 9]. For example, [6] introduces a fusion-prompt pool to complement visual and textual prompts, [8] replays scene-graph prompts to enhance generalization, and [9] applies K-means clustering to define modality-aware prompt centers.
However, most prompt-based CVQA methods adopt a cross-modal prompt isolation approach, constructing and utilizing visual and textual prompts independently without complementary knowledge understanding [6, 17, 4, 9, 10, 11, 13, 12, 8]. As illustrated in Fig. 1(a), these methods select and process prompts separately for vision and language, without mechanisms for cross-modal interaction. This isolated approach is suboptimal for CVQA, where language features naturally tend to be dominant [18]. Injecting such modality-isolated prompts amplifies the existing imbalance by providing the model with additional information from its already-preferred modality, diminishing the ability to integrate cross-modal representations. As a result, performance degrades as the model increasingly relies on one modality.
@@
Figure 1: Comparison between traditional prompt-based CVQA approaches and MM-Prompt. Previous methods rely on isolated prompt pools, letting modality-isolated biases accumulate, whereas MM-Prompt introduces explicit cross‑modal interactions that yield more balanced predictions over sequential tasks.
@@ 2 Related Work
Continual Visual Question Answering. VQA combines visual and language representations to answer questions about images [19, 20], while continual learning builds models that preserve previous knowledge when learning new information [21]. CVQA integrates these approaches by sequentially updating models on new image-question-answer tasks [6, 17, 4, 9, 10, 11, 13, 12, 8]. Approaches to address forgetting in CL include regularization-based[22, 23], architecture-based[24, 25], replay-based[26, 27], and prompt-based methods [10, 11, 12, 17, 13, 9, 4, 6, 8]. While the first three typically require full fine-tuning, prompt-based methods keep backbones frozen and only update lightweight prompts, offering better efficiency. L2P [11], which selects prompts based on query similarity, Dual Prompt [10], which uses general and expert prompts at different layers and CODA [12], which injects weighted combinations of prompt components. Despite their success in general continual learning, prompt methods struggle with CVQA’s unique challenges. CVQA suffers from catastrophic forgetting [28] and modality imbalance [29, 30]. Existing prompt methods select and inject prompts in a modality-specific pattern, providing more information to the model’s preferred modality, typically language [29], which reinforces modality bias and hinders cross-modal integration. As a result, the visual features are further marginalized, and the model’s performance degrades over time.
@@
Pre-trained models in CL. Pretrained models contain rich, generalizable knowledge for downstream tasks [31, 32, 33, 34, 35]. Among CL approaches, prompt-based methods [6, 17, 4, 9, 10, 11, 13, 12, 8, 36] are the most ideal for pretrained models as they preserve capabilities by keeping backbones frozen. While most prompt methods were developed for unimodal settings [10, 11, 12], several recent works explore multimodal prompt learning [6, 17, 4, 8, 9]. [6] introduced fusion prompt pools with interaction matrices to bridge modalities. [8] employs scene graphs as symbolic representations to encourage generalized learning across modalities. [9] uses a two-stage process with K-means clustering to establish modality-specific centers before merging them. Despite these advances, we argue these approaches remain suboptimal for CVQA. As learning progresses, modality-specific differences accumulate within prompts without explicit balancing mechanisms. The selected prompts inevitably carry modality-imbalanced information, which can further exacerbate modality disparity and impair cross-modal integration when injected into the model.
@@ 3.1 Problem Formulation
Following [4], we formulate CVQA as $T$ sequential incremental tasks corresponding to the training data $\{\mathcal{D}_{1},\cdots,\mathcal{D}_{T}\}$. For task $t$, $\mathcal{D}_{t}=\{(x_{i}^{\text{Q}},x_{i}^{\text{V}},y_{i})\}_{i=1}^{n_{t}}$, where $x_{i}^{\text{V}}$ represents input image, $x_{i}^{\text{Q}}$ denotes questions about this image, and $y_{i}$ represents the corresponding true label. During training, models must learn new visual concepts without revisiting prior data, and retain performance on all previous tasks. In this paper, we consider three incremental learning settings: Question Increment (QI), Class Increment (CI), and Dual Increment (DI) [4]. In QI, each task introduces new question types while sharing all object classes across tasks. In CI, each task adds new object classes with shared all question types. DI combines both, in which each task has $S$ subtasks sharing the same visual classes but introducing new question types, while visual classes differ between tasks.
Existing prompt-based methods [6, 17, 4, 9, 10, 11, 13, 12] employ prompt pools, i.e., sets of learnable prompts with associated keys, to store task-specific knowledge. Prompts are used in two stages: first select via query-key matching with input-derived queries and then inject into encoder layers to guide representation learning. However, existing approaches follow a cross-modal prompt isolation strategy, relying on unimodal features for prompt selection and injection, without cross-modal interaction. This reinforces alignment with modality-specific feature distributions, amplifies the dominant modality bias, and hinders the integration of complementary information. To overcome this, we propose MM-Prompt, which introduces two core components: (1) cross-modal prompt query, enabling prompt selection guided by fused modality signals, and (2) cross-modal prompt recovery, facilitating explicit information exchange across modalities. These components are detailed in Fig. 2.
@@ 3.2 Cross-Modal Prompt Query
In CVQA, vision and language offer complementary representations that benefit from joint modeling [19]. Inspired by this, we design a cross-modal querying mechanism that enables each modality to attend to prompts using information from the other. Specifically, we employ an attention (A) module to infuse visual context into question-side queries and linguistic context into image-side queries. For each modality, we maintain a prompt pool consisting of prompts $\mathbf{p}\in\mathcal{P}$ and corresponding keys $\mathbf{k}\in\mathcal{K}$, where $\mathcal{P}$ and $\mathcal{K}$ denote the sets of prompts and keys, respectively, with $\mathbf{p}\in\mathbb{R}^{d}$ and $\mathbf{k}\in\mathbb{R}^{d}$. Let $\mathbf{F}^{\text{Q}}\in\mathbb{R}^{l_{\text{Q}}\times d}$ and $\mathbf{F}^{\text{V}}\in\mathbb{R}^{l_{\text{V}}\times d}$ denote the question and vision features extracted from the $x^{\text{Q}},x^{\text{V}}$, from which we derive the cross-modal queries $\mathbf{q}\in\mathbb{R}^{d}$ with $\mathbf{w}\in\mathbb{R}^{d}$:
$\mathbf{q}^{\text{Q}}=\mathbf{w}^{\text{Q}}\odot\Phi\bigl{(}\text{A}(\mathbf{F}^{\text{Q}},\mathbf{F}^{\text{V}},\mathbf{F}^{\text{V}})+\mathbf{F}^{\text{Q}}\bigr{)},\quad\mathbf{q}^{\text{V}}=\mathbf{w}^{\text{V}}\odot\Phi\bigl{(}\text{A}(\mathbf{F}^{\text{V}},\mathbf{F}^{\text{Q}},\mathbf{F}^{\text{Q}})+\mathbf{F}^{\text{V}}\bigr{)},$ (1)
where $\Phi$ denotes a pooling operation on the feature dimension. To enable cross-modal integration while preserving modality-specific identity, we enhance the query computation in Eq. (1) with two key components. First, the residual term ($+\mathbf{F}^{\text{Q}}/\mathbf{F}^{\text{V}}$) retains original modality features during fusion, preventing dilution by cross-modal signals. Second, the weight modulation terms $\mathbf{w}^{\text{Q}}$ and $\mathbf{w}^{\text{V}}$, initialized uniformly and jointly optimized, adaptively control each feature dimension’s contribution, emphasizing informative components and suppressing noise. Together, these mechanisms ensure that the enriched queries $\mathbf{q}^{\text{Q}}$ and $\mathbf{q}^{\text{V}}$ integrate complementary cross-modal information without compromising modality fidelity, mitigating uni-modal bias and enabling balanced prompt selection.
Figure 2: Detailed MM-Prompt Components for (a) Cross-Modal Prompts Query and (b) Cross-Modal Prompts Recovery from the overall architecture for MM-Prompt. Cross-modal prompts query illustrates our approach where visual (VF) and question (QF) features undergo cross-modal interaction before key matching, creating enriched queries that influence prompt selection across modalities. Cross-modal prompts recovery shows our hierarchical process: identical masks are applied to both modality prompts, followed by intra-modality recovery and inter-modality recovery. This design creates balanced integration pathways that prevent the modality isolation problem in existing prompt-based approaches.
@@ 3.3 Cross-Modal Prompts Recovery
Cross-Modal Masking for Alignment and Dependency. To induce structured interdependence between modalities and encourage explicit cross-modal reasoning, we introduce a shared masking strategy inspired by advances in masked representation learning [37, 38]. Specifically, we apply identical binary masks to both visual and textual prompt sets, thereby creating aligned missing regions across modalities. These masked regions will be reconstructed using complementary information from the other modality, compelling the model to perform genuine cross-modal integration rather than relying on modality-specific priors.
@@
$\mathbf{p}^{\text{Q}}_{\text{intra}}=\text{A}(\mathbf{\hat{p}}^{\text{Q}})+\mathbf{w}_{\text{res}}^{\text{Q}}\odot\mathbf{\tilde{p}}^{\text{V}},\quad\mathbf{p}^{\text{V}}_{\text{intra}}=\text{A}(\mathbf{\hat{p}}^{\text{V}})+\mathbf{w}_{\text{res}}^{\text{V}}\odot\mathbf{\tilde{p}}^{\text{Q}}.$ (5)
@@
$\mathcal{L}_{\text{intra}}=\sum\nolimits_{M\in\{\text{Q},\text{V}\}}||\mathbf{p}^{M}_{\text{intra}}-\mathbf{\tilde{p}}^{M}||_{2}^{2}+||\mathbf{w}_{\text{res}}^{\text{Q}}(\mathbf{w}_{\text{res}}^{\text{Q}})^{T}-\mathbf{I}||_{F}^{2}+||\mathbf{w}_{\text{res}}^{\text{V}}(\mathbf{w}_{\text{res}}^{\text{V}})^{T}-\mathbf{I}||_{F}^{2},$ (6)
where the $||\cdot||_{F}$ represents the Frobenius norm and $\mathbf{I}$ is an identity matrix. By minimizing this loss, the intra-modal recovery establishes a foundation that preserves essential modality-specific patterns while introducing initial cross-modal awareness, creating balanced representations that serve as the basis for deeper integration. However, it still remains insufficient for modeling richer interactions that require explicit cross-modality reasoning. To address this, we introduce a dedicated inter-modal prompt recovery mechanism that operates on the recovered representations to achieve deeper integration and complete the cross-modal reconstruction process.
@@
$\mathbf{p}^{\text{Q}}_{\text{inter}}=\text{A}(\mathbf{p}^{\text{Q}}_{\text{intra}},\mathbf{\tilde{p}}^{\text{V}},\mathbf{\tilde{p}}^{V})+\mathbf{p}^{\text{Q}}_{\text{intra}},\quad\mathbf{p}^{\text{V}}_{\text{inter}}=\text{A}(\mathbf{p}^{\text{V}}_{\text{intra}},\mathbf{\tilde{p}}^{\text{Q}},\mathbf{\tilde{p}}^{\text{Q}})+\mathbf{p}^{\text{V}}_{\text{intra}}.$ (7)
@@
$\mathbf{p}^{\text{Q}}_{\text{final}}=(1-\mathbf{g}^{\text{Q}})\cdot\mathbf{p}^{\text{Q}}_{\text{inter}}+\mathbf{g}^{\text{Q}}\cdot\text{A}(\mathbf{p}^{\text{Q}}_{\text{inter}}),\quad\mathbf{p}^{\text{V}}_{\text{final}}=(1-\mathbf{g}^{\text{V}})\cdot\mathbf{p}^{\text{V}}_{\text{inter}}+\mathbf{g}^{\text{V}}\cdot\text{A}(\mathbf{p}^{\text{V}}_{\text{inter}}),$ (8)
@@
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] represents the concatenation along the feature dimension and $\mathbf{w}_{g}\in\mathbb{R}^{d}$ is a learnable weight matrix that allows the model to adaptively select the areas for enhancement. This design ensures that additional refinement is only performed when supported by strong cross-modal evidence, thus avoiding unnecessary interference and preserving representational sparsity.
To improve inter-modal consistency, we introduce a cross-modal alignment loss that enforces semantic agreement between dual perspectives—how question features are contextualized by vision and vice versa:
$\mathcal{L}_{\text{inter}}=1-\frac{\text{A}(\mathbf{p}^{\text{Q}}_{\text{final}},\mathbf{\tilde{p}}^{\text{V}},\mathbf{\tilde{p}}^{\text{V}})\cdot\text{A}(\mathbf{p}^{\text{V}}_{\text{final}},\mathbf{\tilde{p}}^{\text{Q}},\mathbf{\tilde{p}}^{\text{Q}})}{||\text{A}(\mathbf{p}^{\text{Q}}_{\text{final}},\mathbf{\tilde{p}}^{\text{V}},\mathbf{\tilde{p}}^{\text{V}})||_{2}\cdot||\text{A}(\mathbf{p}^{\text{V}}_{\text{final}},\mathbf{\tilde{p}}^{\text{Q}},\mathbf{\tilde{p}}^{\text{Q}})||_{2}}.$ (10)
By minimizing this loss, we encourage bidirectional semantic coherence and prevent representational drift as the model adapts to new tasks.

### v2

@@ 1 Introduction
Pre-trained models (PTMs) have achieved considerable success in Visual Question Answering (VQA) Lin et al. (2022a); Ravi et al. (2023); Li et al. (2021), a traditional multi-modal task that generates answers to questions based on relevant images. Recently, researchers further extended PTMs to Continual VQA (CVQA) Zhang et al. (2023); Nikandrou et al. (2024); Qian et al. (2023); Lin et al. (2022b); Lei et al. (2023); Cai and Rostami (2024), where new content emerges over time, with approaches developed to address the challenge of catastrophic forgetting. Prompt tuning, a PTM-based method proven effective in CL Wang et al. (2022a, b); Smith et al. (2023); Menabue et al. (2024); Lyu et al. (2021); Sun et al. (2022); Lyu et al. (2023), which first selects prompts based on input representations via similarity, then injects the selected prompts into encoders to guide learning. This approach has shown good performance in CVQA Qian et al. (2023); Zhang et al. (2023); Lei et al. (2023); Cai and Rostami (2024). For example, Qian et al. (2023) introduces a fusion-prompt pool to complement visual and textual prompts, Lei et al. (2023) replays scene-graph prompts to enhance generalization, and Cai and Rostami (2024) applies K-means clustering to define modality-aware prompt centers.
However, most prompt-based CVQA methods adopt a cross-modal prompt isolation approach, constructing and utilizing visual and textual prompts independently without complementary knowledge understanding Qian et al. (2023); Khattak et al. (2023); Zhang et al. (2023); Cai and Rostami (2024); Wang et al. (2022a, b); Menabue et al. (2024); Smith et al. (2023); Lei et al. (2023). As illustrated in Fig. 1(a), these methods select and process prompts separately for vision and language, without mechanisms for cross-modal interaction. This isolated approach is suboptimal for CVQA, where language features naturally tend to be dominant Ramakrishnan et al. (2018). Injecting such modality-isolated prompts amplifies the existing imbalance by providing the model with additional information from its already-preferred modality, diminishing the ability to integrate cross-modal representations. As a result, performance degrades as the model increasingly relies on one modality.
@@
Figure 1: Comparison between traditional prompt-based CVQA approaches and $M$-Prompt. Previous methods rely on isolated prompt pools, letting modality-isolated biases accumulate, whereas MM-Prompt introduces explicit cross‑modal interactions that yield more balanced predictions over sequential tasks.
@@ 2 Related Work
Continual Visual Question Answering. VQA combines visual and language representations to answer questions about images (Goyal et al., 2017; Antol et al., 2015), while continual learning builds models that preserve previous knowledge when learning new information (Thrun, 1995). CVQA integrates these approaches by sequentially updating models on new image-question-answer tasks Qian et al. (2023); Khattak et al. (2023); Zhang et al. (2023); Cai and Rostami (2024); Wang et al. (2022a, b); Menabue et al. (2024); Smith et al. (2023); Lei et al. (2023). Approaches to address forgetting in CL include regularization-based(Zheng et al., 2023; Lao et al., 2023), architecture-based(Yu et al., 2024a; Li et al., 2019), replay-based(Rolnick et al., 2019; Buzzega et al., 2020), and prompt-based methods (Wang et al., 2022a, b; Smith et al., 2023; Khattak et al., 2023; Menabue et al., 2024; Cai and Rostami, 2024; Zhang et al., 2023; Qian et al., 2023; Lei et al., 2023). While the first three typically require full fine-tuning, prompt-based methods keep backbones frozen and only update lightweight prompts, offering better efficiency. L2P (Wang et al., 2022b), which selects prompts based on query similarity, Dual Prompt (Wang et al., 2022a), which uses general and expert prompts at different layers and CODA (Smith et al., 2023), which injects weighted combinations of prompt components. Despite their success in general continual learning, prompt methods struggle with CVQA’s unique challenges. CVQA suffers from catastrophic forgetting (De Lange et al., 2021) and modality imbalance (Yu et al., 2024b; Peng et al., 2022). Existing prompt methods select and inject prompts in a modality-specific pattern, providing more information to the model’s preferred modality, typically language (Yu et al., 2024b), which reinforces modality bias and hinders cross-modal integration. As a result, the visual features are further marginalized, and the model’s performance degrades over time.
@@
Pre-trained Models with Prompt Tuning in CL. Pretrained models contain rich, generalizable knowledge for downstream tasks (Han et al., 2021; Raffel et al., 2020; Tan and Bansal, 2019; Douillard et al., 2022; Radford et al., 2021). Among CL approaches, prompt-based methods Qian et al. (2023); Khattak et al. (2023); Zhang et al. (2023); Cai and Rostami (2024); Wang et al. (2022a, b); Menabue et al. (2024); Smith et al. (2023); Lei et al. (2023); Chang et al. (2024) are the most ideal for pretrained models as they preserve capabilities by keeping backbones frozen. While most prompt methods were developed for unimodal settings (Wang et al., 2022a, b; Smith et al., 2023), several recent works explore multimodal prompt learning (Qian et al., 2023; Khattak et al., 2023; Zhang et al., 2023; Lei et al., 2023; Cai and Rostami, 2024). (Qian et al., 2023) introduced fusion prompt pools with interaction matrices to bridge modalities. (Lei et al., 2023) employs scene graphs as symbolic representations to encourage generalized learning across modalities. (Cai and Rostami, 2024) uses a two-stage process with K-means clustering to establish modality-specific centers before merging them. Despite these advances, we argue these approaches remain suboptimal for CVQA. As learning progresses, modality-specific differences accumulate within prompts without explicit balancing mechanisms. The selected prompts inevitably carry modality-imbalanced information, which can further exacerbate modality disparity and impair cross-modal integration when injected into the model.
Cross Modal Interaction A number of vision–language models have explored explicit cross-modal alignment and fusion strategies in static, single-task settings (Nagrani et al., 2021; Wu et al., 2022; Liu et al., 2019). Attention Bottlenecks (Nagrani et al., 2021) align modalities by routing inter-modal information through a small set of shared bottleneck tokens, with early independent processing followed by mid-level fusion. DIFNet (Wu et al., 2022) aligns complementary visual streams using Iterative Independent Layer Normalization, preserving modality-specific cues via shared weights and separate normalization. MIA (Liu et al., 2019) uses mutual iterative attention, alternately attending between visual regions and textual concepts before fusing the aligned features. While effective for joint learning, these designs are suboptimal for Continual VQA. Their fusion mechanisms require continual fine-tuning, which overwrites knowledge from previous tasks and makes them highly prone to catastrophic forgetting. In addition, their static fusion strategies lack the adaptability needed to mitigate the modality imbalance that often emerges and accumulated across sequential tasks in CVQA.
@@ 3.1 Problem Formulation
Following Zhang et al. (2023), we formulate CVQA as $T$ sequential incremental tasks corresponding to the training data $\{\mathcal{D}_{1},\cdots,\mathcal{D}_{T}\}$. For task $t$, $\mathcal{D}_{t}=\{(x_{i}^{\text{Q}},x_{i}^{\text{V}},y_{i})\}_{i=1}^{n_{t}}$, where $x_{i}^{\text{V}}$ represents input image, $x_{i}^{\text{Q}}$ denotes questions about this image, and $y_{i}$ represents the corresponding true label. During training, models must learn new visual concepts without revisiting prior data, and retain performance on all previous tasks. In this paper, we consider three incremental learning settings: Question Increment (QI), Class Increment (CI), and Dual Increment (DI) (Zhang et al., 2023). In QI, each task introduces new question types while sharing all object classes across tasks. In CI, each task adds new object classes with shared all question types. DI combines both, in which each task has $S$ subtasks sharing the same visual classes but introducing new question types, while visual classes differ between tasks.
Existing prompt-based methods Qian et al. (2023); Khattak et al. (2023); Zhang et al. (2023); Cai and Rostami (2024); Wang et al. (2022a, b); Menabue et al. (2024); Smith et al. (2023) employ prompt pools, i.e., sets of learnable prompts with associated keys, to store task-specific knowledge. Prompts are used in two stages: first select via query-key matching with input-derived queries and then inject into encoder layers to guide representation learning. However, existing approaches follow a cross-modal prompt isolation strategy, relying on unimodal features for prompt selection and injection, without cross-modal interaction. This reinforces alignment with modality-specific feature distributions, amplifies the dominant modality bias, and hinders the integration of complementary information. To overcome this, we propose MM-Prompt, which introduces two core components: (1) cross-modal prompt query, enabling prompt selection guided by fused modality signals, and (2) cross-modal prompt recovery, facilitating explicit information exchange across modalities. These components are detailed in Fig. 2.
@@ 3.2 Cross-Modal Prompt Query
In CVQA, vision and language offer complementary representations that benefit from joint modeling (Goyal et al., 2017). Inspired by this, we design a cross-modal querying mechanism that enables each modality to attend to prompts using information from the other. Specifically, we employ an attention (A) module to infuse visual context into question-side queries and linguistic context into image-side queries. For each modality, we maintain a prompt pool consisting of prompts $\mathbf{p}\in\mathcal{P}$ and corresponding keys $\mathbf{k}\in\mathcal{K}$, where $\mathcal{P}$ and $\mathcal{K}$ denote the sets of prompts and keys, respectively, with $\mathbf{p}\in\mathbb{R}^{d}$ and $\mathbf{k}\in\mathbb{R}^{d}$. Let $\mathbf{F}^{\text{Q}}\in\mathbb{R}^{l_{\text{Q}}\times d}$ and $\mathbf{F}^{\text{V}}\in\mathbb{R}^{l_{\text{V}}\times d}$ denote the question and vision features extracted from the $x^{\text{Q}},x^{\text{V}}$, from which we derive the cross-modal queries $\mathbf{q}\in\mathbb{R}^{d}$ with $\mathbf{w}\in\mathbb{R}^{d}$:
$\mathbf{q}^{\text{Q}}=\mathbf{w}^{\text{Q}}\odot\Phi\bigl{(}\text{A}_{\text{query}}^{\text{Q}}(\mathbf{F}^{\text{Q}},\mathbf{F}^{\text{V}},\mathbf{F}^{\text{V}})+\mathbf{F}^{\text{Q}}\bigr{)},\quad\mathbf{q}^{\text{V}}=\mathbf{w}^{\text{V}}\odot\Phi\bigl{(}\text{A}_{\text{query}}^{\text{V}}(\mathbf{F}^{\text{V}},\mathbf{F}^{\text{Q}},\mathbf{F}^{\text{Q}})+\mathbf{F}^{\text{V}}\bigr{)},$ (1)
where $\Phi$ denotes a pooling operation on the feature dimension. To enable cross-modal integration while preserving modality-specific identity, we enhance the query computation in Eq. (1) with two key components. First, the residual term ($\mathbf{F}^{\text{Q}}/\mathbf{F}^{\text{V}}$) retains original modality features during fusion, preventing dilution by cross-modal signals. Second, the weight modulation terms $\mathbf{w}^{\text{Q}}$ and $\mathbf{w}^{\text{V}}$, initialized uniformly and jointly optimized, adaptively control each feature dimension’s contribution, emphasizing informative components and suppressing noise. Together, these mechanisms ensure that the enriched queries $\mathbf{q}^{\text{Q}}$ and $\mathbf{q}^{\text{V}}$ integrate complementary cross-modal information without compromising modality fidelity, mitigating uni-modal bias and enabling balanced prompt selection.
Figure 2: Detailed $M$-Prompt Components for (a) Cross-Modal Prompts Query and (b) Cross-Modal Prompts Recovery from the overall architecture for $M$-Prompt. Cross-modal prompts query illustrates our approach where visual (VF) and question (QF) features undergo cross-modal interaction before key matching, creating enriched queries that influence prompt selection across modalities. Cross-modal prompts recovery shows our hierarchical process: identical masks are applied to both modality prompts, followed by intra-modality recovery and inter-modality recovery. This design creates balanced integration pathways that prevent the modality isolation problem in existing prompt-based approaches.
@@ 3.3 Cross-Modal Prompts Recovery
Cross-Modal Masking for Alignment and Dependency. To induce structured interdependence between modalities and encourage explicit cross-modal reasoning, we introduce a shared masking strategy inspired by advances in masked representation learning (He et al., 2022; Zhang et al., 2022). Specifically, we apply identical binary masks to both visual and textual prompt sets, thereby creating aligned missing regions across modalities. These masked regions will be reconstructed using complementary information from the other modality, compelling the model to perform genuine cross-modal integration rather than relying on modality-specific priors.
@@
$\mathbf{p}^{\text{Q}}_{\text{intra}}=\text{A}_{\text{intra}}^{\text{Q}}(\mathbf{\hat{p}}^{\text{Q}})+\mathbf{w}_{\text{res}}^{\text{Q}}\odot\mathbf{\tilde{p}}^{\text{V}},\quad\mathbf{p}^{\text{V}}_{\text{intra}}=\text{A}_{\text{intra}}^{\text{V}}(\mathbf{\hat{p}}^{\text{V}})+\mathbf{w}_{\text{res}}^{\text{V}}\odot\mathbf{\tilde{p}}^{\text{Q}}.$ (5)
@@
[…]
where the $||\cdot||_{F}$ represents the Frobenius norm and $\mathbf{I}$ is an identity matrix. First term preserves the modality-specific information by keeping the recovered prompt close to the original selected prompt, while second and third term enforce orthogonality on the residual weights to limit redundant or overly correlated dimensions (Bansal et al., 2018). By minimizing this loss, the intra-modal recovery establishes a foundation that preserves essential modality-specific patterns while introducing initial cross-modal awareness, creating balanced representations that serve as the basis for deeper integration. However, it still remains insufficient for modeling richer interactions that require explicit cross-modality reasoning. To address this, we introduce a dedicated inter-modal prompt recovery mechanism that operates on the recovered representations to achieve deeper integration and complete the cross-modal reconstruction process.
@@
$\mathbf{p}^{\text{Q}}_{\text{inter}}=\text{A}_{\text{inter}}^{\text{Q}}(\mathbf{p}^{\text{Q}}_{\text{intra}},\mathbf{\tilde{p}}^{\text{V}},\mathbf{\tilde{p}}^{V})+\mathbf{p}^{\text{Q}}_{\text{intra}},\quad\mathbf{p}^{\text{V}}_{\text{inter}}=\text{A}_{\text{inter}}^{\text{V}}(\mathbf{p}^{\text{V}}_{\text{intra}},\mathbf{\tilde{p}}^{\text{Q}},\mathbf{\tilde{p}}^{\text{Q}})+\mathbf{p}^{\text{V}}_{\text{intra}}.$ (7)
@@
$\mathbf{p}^{\text{Q}}_{\text{final}}=(1-\mathbf{g}^{\text{Q}})\cdot\mathbf{p}^{\text{Q}}_{\text{inter}}+\mathbf{g}^{\text{Q}}\cdot\text{A}_{\text{gate}}^{\text{Q}}(\mathbf{p}^{\text{Q}}_{\text{inter}}),\quad\mathbf{p}^{\text{V}}_{\text{final}}=(1-\mathbf{g}^{\text{V}})\cdot\mathbf{p}^{\text{V}}_{\text{inter}}+\mathbf{g}^{\text{V}}\cdot\text{A}_{\text{gate}}^{\text{V}}(\mathbf{p}^{\text{V}}_{\text{inter}}),$ (8)
@@
where $\cdot$$\cdot$] represents the concatenation along the feature dimension and $\mathbf{w}_{g}\in\mathbb{R}^{d}$ is a learnable weight matrix that allows the model to adaptively select the areas for enhancement. This design ensures that additional refinement is only performed when supported by strong cross-modal evidence, thus avoiding unnecessary interference and preserving representational sparsity.
To improve inter-modal consistency, we introduce a cross-modal alignment loss that enforces semantic agreement between dual perspectives—how question features are contextualized by vision and vice verse (Numerator):
[…]
where Denominator normalizes both scores to focus on directional agreement rather than magnitude. By minimizing this loss, we encourage bidirectional semantic coherence and prevent representational drift as the model adapts to new tasks.